is i had to solve for y what would x+y=-3
Answer by unlockmath(1688)

Hello,
To solve for y in the equation x + y = -3, we need to get y by itself on one side of the equal sign. Here's how you do it.
You subtract x on both sides of the equal sign. (remember,
whatever you do on one side of the equal sign you must do on
one the other side.)
x - x + y = -3 -x (You see, x - x is zero therefore leaves y all alone.)
y = -3 - x (Here's your answer)
RJ Toftness
www.math-unlock.com